OpenStudy (anonymous):

Hello everyone can I ask help regarding visual basic. Basic Data types in VB, DimVariable,Declaration,arithmetic operators. I want to ask help from you guys what code of these. Please help me and give me examples because we have a report on Monday then we have to demonstrate step by step all of these.

OpenStudy (anonymous):

everyone help me

OpenStudy (anonymous):

You can declare variables like: Dim x as Number Dim y as Number Dim sum as Number Dim product as Number Dim z as String and then use them as arithmetic expressions like so x = 10 y = 10 z = "hello" sum = x + y product = x * y and so one you can check this link for more information: http://www.csidata.com/custserv/onlinehelp/vbsdocs/vbs13.htm